Draw the major organic product for the reaction. Draw the product. Select Draw Rings More HgSO4, H20 H2SO4 = H. ОН
Answer #1

The solution is given as follows :- o-soo T ush My SoysaaH → ellito so , H9-0-500 Hg-o-soo -c =é-H POH2 -cac-H - H+ -C=CH₂ SD